# Erratic non-Hermitian skin localization

### v1

@@ Abstract
Introduction. Anderson localization R1 ; R2 ; R3 ; R4 and the non-Hermitian skin effect (NHSE) R5 ; R6 ; R7 ; R8 ; R9 ; R10 ; R11 ; R11b ; R12 ; R13 ; R14 ; R15 ; R16 ; R17 ; R17b ; R18 ; R19 ; R20 ; R21 ; R21b ; R21c ; R22 ; R23 ; R24 ; R25 ; R26 ; R27 ; R28 ; R28b ; R29 ; R30 ; R31 represent two fundamental wave localization phenomena, arising from distinct mechanisms: wave interference in disordered systems and the intrinsic point-gap topology of non-Hermitian (NH) systems, respectively. The Hatano-Nelson model has long served as a cornerstone in the study of NH systems, first introduced in seminal works R32 ; R33 ; R34 . Earlier research on this model uncovered a competition between Anderson localization and delocalization induced by an imaginary gauge field R32 ; R33 ; R34 ; R35 ; R36 ; R37 ; R38 ; R39 ; R40 ; R41 ; R42 ; R43 . Recent advances in non-Hermitian and topological physics R20 ; R43a ; R43b ; R43c ; R43d ; R43e have unveiled a rich variety of phenomena beyond these initial discoveries, highlighting the exceptional sensitivity of NH systems to boundary conditions, as demonstrated by the NHSE (see, e.g., R5 ; R7 ; R20 ; R22 ; R23 ; R26 ; R27 ; R28 ; R31 and references therein). This phenomenon, characterized by the accumulation of a large number of eigenstates near the system boundaries R5 , defies the conventional Bloch band theory and challenges the traditional bulk-boundary correspondence. To reconcile this discrepancy, the generalized Brillouin zone theory R5 ; R8 ; R44 in one dimensional systems, and amoeba formulation of non-Bloch band theory in higher dimensions R44bis , have been introduced. The NHSE can persist in disordered systems with broken translational invariance R55 ; R56 ; R56b ; R57 ; R58 ; R59 ; R59b ; R59c . The interplay between non-Hermiticity and spatial inhomogeneities –such as domain walls, disorder, dislocations or impurities– gives rise to intriguing localization phenomena R27 , including topological phase transitions R44a ; R44b ; R44c ; R44d , impurity-induced topological bound states R17 ; R45 ; R46 ; R47 , scale-free localization R48 ; R48b ; R49 ; R50 ; R51 ; R52 ; R53 ; R54 , dislocation NHSE RD1 ; RD2 ; RD3 , the inner NHSE RD4 and unusual form of NH transport R54a ; R54b ; R54c ; R54d ; R54e .

### v2

@@ Abstract
Introduction. Anderson localization R1 ; R2 ; R3 ; R4 and the non-Hermitian skin effect (NHSE) R5 ; R6 ; R7 ; R8 ; R9 ; R10 ; R11 ; R11b ; R11c ; R12 ; R13 ; R14 ; R15 ; R16 ; R17 ; R17b ; R18 ; R19 ; R20 ; R21 ; R21b ; R21c ; R22 ; R23 ; R24 ; R25 ; R26 ; R27 ; R28 ; R28b ; R29 ; R30 ; R31 represent two fundamental wave localization phenomena, arising from distinct mechanisms: wave interference in disordered systems and the intrinsic point-gap topology of non-Hermitian (NH) systems, respectively. The Hatano-Nelson model has long served as a cornerstone in the study of NH systems, first introduced in seminal works R32 ; R33 ; R34 . Earlier research on this model uncovered a competition between Anderson localization and delocalization induced by an imaginary gauge field R32 ; R33 ; R34 ; R35 ; R36 ; R37 ; R38 ; R39 ; R40 ; R41 ; R42 ; R43 . Recent advances in non-Hermitian and topological physics R20 ; R43a ; R43b ; R43c ; R43d ; R43e have unveiled a rich variety of phenomena beyond these initial discoveries, highlighting the exceptional sensitivity of NH systems to boundary conditions, as demonstrated by the NHSE (see, e.g., R5 ; R7 ; R20 ; R22 ; R23 ; R26 ; R27 ; R28 ; R31 and references therein). This phenomenon, characterized by the accumulation of a large number of eigenstates near the system boundaries R5 , defies the conventional Bloch band theory and challenges the traditional bulk-boundary correspondence. To reconcile this discrepancy, the generalized Brillouin zone theory R5 ; R8 ; R44 in one dimensional systems, and amoeba formulation of non-Bloch band theory in higher dimensions R44bis , have been introduced. The NHSE can persist in disordered systems with broken translational invariance R55 ; R56 ; R56b ; R57 ; R58 ; R59 ; R59b ; R59c . The interplay between non-Hermiticity and spatial inhomogeneities –such as domain walls, disorder, dislocations or impurities– gives rise to intriguing localization phenomena R27 , including topological phase transitions R44a ; R44b ; R44c ; R44d , impurity-induced topological bound states R17 ; R45 ; R46 ; R47 , scale-free localization R48 ; R48b ; R49 ; R50 ; R51 ; R52 ; R53 ; R54 , dislocation NHSE RD1 ; RD2 ; RD3 , the inner NHSE RD4 and unusual form of NH transport R54a ; R54b ; R54c ; R54d ; R54e .